OpenStudy (jojo12g):

Translate the sentence into an equation. Seven times the difference of a number and 3 is 35. A. 7k – 3 = 35 B. k – 3 ∙ 7 = 35 C. 7 ∙ (k – 3) = 35 D. 7 ∙ (k + 3) = 35

OpenStudy (anonymous):

most people would start it from the left of the sentence to the right, but than it be incorrect so you should always start in the middle so it said the different between number and 3 so it (k-3)

OpenStudy (anonymous):

is which mean in equal difference means subtract and times is well duh * so equal to 35 and it times to 7 so it be C ;)

OpenStudy (mathstudent55):

\(\sf \color{blue}{Seven ~times} \color{red}{~the ~difference ~of ~a ~number ~and ~3} \color{green}{~is} \color{purple}{~35}.\) \(\sf \color{blue}{~~~~~~~~~7~\cdot~~~~} ~~~~~~~~~~~~~~~~~~~~~~~~ \color{red}{(k - 3)} ~~~~~~~~~~~~~~~~~~~\color{green}{=} \color{purple}{35}\) \(7 \cdot (k - 3) = 35\)